OpenStudy (anonymous):

what is (5π)/4 radians in degrees?

Directrix (directrix):

Pi radians is equivalent to 180 degrees. Set up a proportion: (5π)/4 / π = x / 180 5/4 = x /180 4x = 5*180 x = ? degrees

OpenStudy (anonymous):

225

Directrix (directrix):

Correct. 225 degrees.